Ticket Sales
Suppose you are an event coordinator for a large performance theater. One of the hottest new Broadway musicals has started to tour, and your city is the first stop on the tour. You need to supply information about projected ticket sales to the box office manager. The box office manager uses this information to anticipate staffing needs until the tickets sell out. You provide the manager with a quadratic equation that models the expected number of ticket sales for each day x. (x=1 is the day tickets go on sale). Tickets = -0.2x^2+12x+11
a.Does the graph of this equation open up or down? How did you determine this?
b.Describe what happens to the tickets sales as time passes?
c.Use the quadratic equation to determine the last day that tickets will be sold.(Note write your answer in terms of the number of days after ticket sales begin.)
d.Will tickets peak or be at a low during the middle of the sale?How do you know?
e.After how many days will the peak or low occur?
f.How many tickets will be sold on the day when the peak or low occurs?
g.What is the point of the vertex?How does this number relate to your answers in parts e and f?
h.How many solutions are there to the equation -2x^2+12x+11=0? How do you know?
i.What do the solutions represent? Is there a solution that does not make sense? If so, in what ways does the solution not make sense?

Answer by stanbon(75887)

You provide the manager with a quadratic equation that models the expected number of ticket sales for each day x. (x=1 is the day tickets go on sale). Tickets = -0.2x^2+12x+11
a.Does the graph of this equation open up or down? How did you determine this?
Ans: Down
---------------------
b.Describe what happens to the tickets sales as time passes?
Ans Up then down.
---------------------
c.Use the quadratic equation to determine the last day that tickets will be sold.(Note write your answer in terms of the number of days after ticket sales begin.)
Ans: Solve -0.2x^2 + 12x + 11 = 0
-----------------------------
d.Will tickets peak or be at a low during the middle of the sale?How do you know?
Ans: Peak
-----------------------------------
e.After how many days will the peak or low occur?
Ans: Peak occurs when x = -b/2a = -12/(2*-0.2) = 30
--------------
f.How many tickets will be sold on the day when the peak or low occurs?
Ans: Find f(30)
----------------
g.What is the point of the vertex?How does this number relate to your answers in parts e and f?
Ans: (30,f(3))
----------------------------
h.How many solutions are there to the equation -2x^2+12x+11=0? How do you know?
Ans: Check b^2-4ac = 12^2-4*-2*11 > 0 so two Real Number solutions.
-------------------------------
i.What do the solutions represent? Is there a solution that does not make sense? If so, in what ways does the solution not make sense?
Ans: The number of the day when there are zero sales.
====================================================
Cheers,
stan H.